# HiFo-Prompt: Prompting with Hindsight and Foresight for LLM-based Automatic Heuristic Design

### v1

@@ C.4 Bayesian Optimization
Appendix D Algorithm Details
This section provides detailed pseudocode for the core components of the HiFo-Prompt framework. We first present the main algorithm, followed by the specific implementations for the survival selection and parent selection mechanisms.
D.1 Main Framework Algorithm
Algorithm 1 outlines the complete, end-to-end procedure for the HiFo-Prompt framework. It details the main evolutionary loop, starting from population initialization and proceeding through iterative generations. Each generation consists of four primary phases: (1) Foresight and Hindsight, where the Evolutionary Navigator and Insight Pool generate the control vector; (2) Construction, where the LLM generates new heuristics under the guidance of the control vector; (3) Evaluation & Feedback, where offspring are evaluated and the Insight Pool learns from the results; and (4) Selection, where the next generation’s population is formed.
Algorithm 1 HiFo-Prompt
Input: Problem definition $\mathcal{F}$, optional seed algorithms $\mathcal{P}_{\text{seed}}$
Parameter: Population size $N_{\text{pop}}$, number of generations $G_{\max}$, set of evolutionary foundation prompt strategies $\mathcal{O}$ with weights $\mathcal{W}$, number of parents $m$
Output: The final population of evolved algorithms $\mathcal{P}$
1: Initialize Population $\mathcal{P}$:
2:if $\mathcal{P}_{\text{seed}}$ is provided then
3:  $\mathcal{P}\leftarrow$ Evaluate and populate from $\mathcal{P}_{\text{seed}}$
4:else
5:  $\mathcal{P}\leftarrow\emptyset$
6:  while $|\mathcal{P}|<N_{\text{pop}}$ do
7:   $\text{Ind}_{\text{new}}\leftarrow$ Generate an initial algorithm via LLM
8:   Evaluate fitness of $\text{Ind}_{\text{new}}$ using $\mathcal{F}$
9:   Add valid $\text{Ind}_{\text{new}}$ to $\mathcal{P}$
10:  end while
11:end if
12: Initialize Insight Pool $\mathcal{M}_{\text{insight}}$ with a set of default design principles
13: Initialize Evolutionary Navigator $\mathcal{C}_{\text{meta}}$
14:
15:for $g=1$ to $G_{\max}$ do
16:  Update $\mathcal{C}_{\text{meta}}$ with current population state
17:  if a condition for wisdom extraction is met then
18:   $\mathcal{P}_{\text{top}}\leftarrow$ Select top-performing individuals from $\mathcal{P}$
19:   $\text{new\_tips}\leftarrow$ Prompt LLM to extract generic design principles from $\mathcal{P}_{\text{top}}$
20:   Add new_tips to $\mathcal{M}_{\text{Insight}}$
21:  end if
22:
23:  for each foundation prompt strategy $o\in\mathcal{O}$ do
24:   if $\text{random}()<\mathcal{W}[o]$ then
[…]
26:    $\text{insights}\leftarrow\mathcal{M}_{\text{Insight}}$.get_insights() {Get inspiring principles}
27:    $\mathcal{P}_{\text{parents}}\leftarrow$ Select $m$ parents from $\mathcal{P}$
28:    $\text{Ind}_{\text{new}}\leftarrow$ Generate offspring via LLM using $o$, $\mathcal{P}_{\text{parents}}$, guidance, and insights
29:    Evaluate fitness of $\text{Ind}_{\text{new}}$ using $\mathcal{F}$
30:    if $\text{Ind}_{\text{new}}$ is valid and not duplicated then
31:     $\text{effectiveness}\leftarrow$ Calculate effectiveness of tips based on $\text{Ind}_{\text{new}}$’s performance
32:     $\mathcal{M}_{\text{Insight}}$.update_insight_stats(insights, effectiveness) {Feedback loop}
33:     Add $\text{Ind}_{\text{new}}$ to $\mathcal{P}$
34:    end if
35:   end if
36:  end for
37:  $\mathcal{P}\leftarrow$ SurvivalSelection($\mathcal{P}$, $N_{\text{pop}}$) {Manage population size}
38:end for
39:return Best individuals from $\mathcal{P}$
D.2 Survival Selection Mechanism
To maintain a constant population size across generations, we employ a survival selection mechanism after new offspring have been generated and evaluated. Algorithm 2 details this procedure. It implements a greedy, elitist strategy that first filters out any invalid or unsuccessfully evaluated individuals. It then ensures uniqueness based on objective values to maintain diversity in the performance space. Finally, it sorts the unique, valid individuals by their performance and truncates the population to the target size $N_{target}$, ensuring that only the highest-performing heuristics are retained for the next generation.
Algorithm 2 Greedy Population Management
Input: A population of individuals $P$, Target population size $N_{target}$
Output: A new population $P_{new}$ with size at most $N_{target}$
[…]
2: Initialize an empty list $P_{unique}$
3: Initialize an empty set of seen objectives $O_{seen}$
4:for each individual $i\in P_{valid}$ do
5:  if $i.objective\notin O_{seen}$ then
6:   Add $i$ to $P_{unique}$
7:   Add $i.objective$ to $O_{seen}$
8:  end if
9:end for
[…]
11: Sort $P_{unique}$ in ascending order based on objective values.
[…]
13:return $P_{new}$
D.3 Parent Selection Mechanism
For the generative operators that require parent heuristics (e.g., crossover and mutation), a parent selection mechanism is used to choose individuals from the current population. Algorithm 3 describes the rank-based selection method used in our framework. This approach assigns a selection probability to each individual that is inversely proportional to its performance rank. This ensures that higher-performing individuals are more likely to be selected as parents, creating a strong selection pressure that guides the search towards promising regions of the heuristic space, while still allowing lower-ranked individuals a chance to contribute.
Algorithm 3 Rank-Based Parent Selection
Input: Sorted population $P$ (best to worst), Number of parents to select $m$
Output: A list of selected parents $P_{parents}$
1: Let $N\leftarrow$ size of $P$
2: Initialize an empty list for probabilities $W\leftarrow\{\}$
3:for $k=0$ to N1N-1 do
4:  Calculate probability $w_{k}\propto 1/(k+1+N)$ {Rank-based weighting}
5:  Append $w_{k}$ to $W$
6:end for
7: Select $m$ individuals from $P$ with replacement, using probabilities $W$, to form $P_{parents}$.
8:return $P_{parents}$
Appendix E Limitation and Future Work
E.1 Limitation

### v2

@@ C.4 Bayesian Optimization
C.5 Comparative Results
Convergence Analysis.
We compared the progression of objective function values during the evolutionary process of our method and EoH on both the TSP and Online BPP. Figure 3 presents the convergence analysis curves. The figure demonstrates that our method converges more rapidly while requiring fewer individuals in the population, indicating higher efficiency. For a detailed breakdown of computational costs, including runtime and token consumption comparisons, please refer to Appendix C.8.
Figure 3: A Key Example of Convergence Analysis for TSP and Online BPP.
Results with Different LLMs.
Furthermore, we evaluated it across a diverse set of LLMs. To ensure fairness and consistency, we adopted the evaluation protocol established in MCTS-AHD  [67]. Specifically, we selected two representative problem scales for each task to validate scalability, and each experiment was conducted across three independent runs, with average values reported to mitigate stochasticity. The detailed comparative results are presented in Table 11.
The results indicate that our framework exhibits strong adaptability, consistently discovering effective heuristics across diverse LLMs. While Qwen-2.5-max yields the leading metrics , we observe that performance does not follow the generic reasoning benchmarks of these models. Instead, success appears to rely on the synergy between the framework’s guidance mechanisms and the specific model’s characteristics (e.g., instruction following). Even smaller models like GPT-4o-mini show competitive results, highlighting that optimizing the coordination between the LLM-based AHD and the underlying model remains a promising and open direction for future research.
Table 11: Results with different LLMs.
Problems TSP-construction TSP-GLS
Scale N=50 N=100 N=100 N=200
Qwen-2.5-max 6.625% 8.582% 0.015% 0.382%
GPT-4o-mini 7.058% 9.579% 0.021% 0.433%
DeepSeek-v3 7.717% 10.455% 0.041% 0.480%
DeepSeek-r1 10.412% 12.638% 0.074% 0.589%
Claude3.5-sonnet 8.276% 11.087% 0.061% 0.454%
Problems Online BPP FSSP-GLS
Scale N=1k, C=100 N=10k, C=300 n=100, m=10 n=200, m=20
Qwen-2.5-max 2.188% 0.391% 0.128% 0.710%
GPT-4o-mini 2.287% 0.492% 0.137% 0.787%
DeepSeek-v3 2.238% 0.978% 0.185% 0.916%
DeepSeek-r1 2.984% 1.487% 0.313% 1.485%
Claude3.5-sonnet 2.089% 1.487% 0.179% 1.256%
C.6 Parameter Sensitivity Analysis
To rigorously quantify the impact of key hyperparameters on the performance of HiFo-Prompt, we conducted a comprehensive ablation study involving eight core parameters. All sensitivity experiments were conducted on the TSP-50 construction task using the training dataset. To ensure statistical reliability, each configuration was executed across 3 independent runs, with results reported as the average objective function value. Unless otherwise specified, the framework operated under a default configuration of population size $N_{\text{pop}}=8$ and maximum generations $G_{\text{max}}=8$. The results confirm that the framework consistently converges to high-quality solutions across a broad spectrum of configurations, demonstrating that our settings are both effective and robust without relying on instance-specific fine-tuning. The results, visualized in Figure 4 and detailed in Table 12, not only justify our default configurations but also reveal the underlying mechanics of the framework. Collectively, these experiments demonstrate that HiFo-Prompt maintains consistent performance across a reasonable range of parameter settings, confirming the robustness and reproducibility of the proposed method.
Recency Bonus Magnitude $(\tau_{r})$
The role of this parameter is grounded in the formulation of the recency bonus function $B_{r}$, designed to capture the temporal dynamics of the search process. Formally defined as:
$B_{r}(t,t_{i}^{\text{last}})=\begin{cases}\tau_{r}&\text{if }t-t_{i}^{\text{last}}\leq T_{w}\\ 0&\text{otherwise}\end{cases}$ (24)
where $t$ denotes the current generation index, $t_{i}^{\text{last}}$ records the last generation in which insight $i$ was successfully applied, and $T_{w}$ represents the sliding window size (set to $T_{w}=2$). Here, $\tau_{r}$ controls the magnitude of the reward assigned to recently validated insights. Theoretically, this mechanism addresses the non-stationary nature of evolutionary search—an insight highly effective during early exploration may become obsolete during later exploitation. This design aligns with the principles of Sliding-Window Upper Confidence Bound (SW-UCB) [18] algorithms used in non-stationary Multi-Armed Bandits. By restricting the bonus to a narrow temporal window $T_{w}$, the framework implicitly discounts outdated information, ensuring that the selection process prioritizes insights that are not just historically successful, but contextually relevant to the current state of the population.
The sensitivity analysis presented in Table 12 validates this design choice. Specifically, when this short-term memory is deactivated ($\tau_{r}=0$), the algorithm inherently lacks the incentive to maintain search continuity. Consequently, the search explores widely but inefficiently, resulting in degraded performance. However, it is crucial to note that simply increasing the reward is not monotonically beneficial. As observed in the table, setting the reward magnitude too high ($\tau_{r}=0.4$) leads to a deterioration in algorithm quality. This outcome suggests that excessive recency bias induces myopic behavior, wherein the algorithm becomes overly focused on the specific traits of the most recent iteration. Such short-sighted focus ultimately overshadows the robust, long-term effectiveness signals ($E_{i}$) accumulated over time. In contrast, introducing a moderate reward ($\tau_{r}=0.2$) achieves consistently strong performance. This finding confirms that explicitly rewarding recent successes allows the LLM to effectively build upon immediate gains, thus guiding the population through the shifting fitness landscape without becoming distracted by transient fluctuations.
Table 12: Ablation analysis of recency bonus magnitude ($\tau_{r}$), usage penalty weight ($w_{u}$) and decay rate ($R_{\text{decay}}$).
$w_{u}$ run1 run2 run3 avg.
0 5.958 5.989 5.978 5.975
0.05 5.894 5.901 5.907 5.901
0.1 (default) 5.885 5.838 5.873 5.866
0.3 5.921 5.907 5.922 5.917
$\tau_{r}$ run1 run2 run3 avg.
0 5.954 5.972 5.979 5.968
0.1 5.920 5.925 5.901 5.915
0.2 (default) 5.885 5.838 5.873 5.866
0.4 5.937 5.945 5.933 5.938
$R_{\text{decay}}$ run1 run2 run3 avg.
0 5.998 6.097 6.075 6.057
0.005 5.899 5.908 5.906 5.904
0.01 (default) 5.885 5.838 5.873 5.866
0.04 5.985 5.956 5.967 5.969
Usage Penalty Weight ($w_{u}$).
This parameter regulates the exploration pressure within the Insight Pool by modulating the usage cost term, defined as $w_{u}\log(N_{i}(t)+1)$, where $N_{i}(t)$ represents the retrieval count of insight $i$. The choice of a logarithmic growth function is deliberate: it ensures that the penalty accumulates rapidly for the first few uses to prevent immediate overfitting, but saturates quickly thereafter. This design enables the score of a frequently used insight to stabilize rather than suffer a precipitous decline, thereby avoiding sharp priority reversals or oscillatory behavior while granting other high-potential candidates a fair chance. The sensitivity analysis in Table 12 confirms the necessity of this balanced regulation. When the penalty is removed ($w_{u}=0$), the framework exhibits a significant performance drop. This indicates a phenomenon of knowledge collapse, where the algorithm over-exploits a few early dominant insights, effectively depriving the search of diversity. Conversely, an aggressive penalty ($w_{u}=0.3$) yields suboptimal results by discarding valid insights prematurely before their full potential is realized. The default value of $w_{u}=0.1$ demonstrates robust efficacy, showing that a gentle, saturating penalty is sufficient to maintain a healthy turnover of valid ideas without suppressing fundamental design principles.
Decay Rate ($R_{\text{decay}}$).
This parameter governs the temporal lifecycle of insights, controlling the rate at which older knowledge becomes obsolete. By applying a linear decay to the eviction score, $R_{\text{decay}}$ enforces a mechanism of knowledge turnover, preventing the Insight Pool from becoming a static archive of legacy strategies that, while successful in early generations, may no longer be competitive. The sensitivity analysis in Table 12 reveals that this dynamic maintenance is critical for sustained performance. When the decay mechanism is disabled ($R_{\text{decay}}=0$), the algorithm yields the worst results among all tested configurations. This confirms that a static pool eventually becomes saturated with stale information, effectively locking out superior novel insights due to capacity constraints. Conversely, setting the rate too high ($R_{\text{decay}}=0.04$) causes the system to rapidly forget valid principles, disrupting the accumulation of algorithmic knowledge. The default rate of $R_{\text{decay}}=0.01$ demonstrates robust performance, continuously purging outdated or inactive insights to ensure the pool remains relevant and valid, thereby staying aligned with the evolving needs of the population.
Diversity Threshold ($\delta_{p}$).
This parameter establishes the minimum acceptable level of phenotypic diversity, acting as a sensitivity trigger for the Evolutionary Navigator. As defined in Eq. 7, diversity is computed via exact string matching of the generated algorithmic descriptions. This rigorous design choice is motivated by the fact that our standardized prompt template eliminates trivial syntactic noise (e.g., variable renaming), ensuring that any remaining lexical difference reflects a deliberate structural modification by the LLM. Moreover, exact matching captures subtle but critical logical alterations—such as dynamic vs. static—that embedding-based metrics might obscure. The threshold $\delta_{p}$ dictates when the system perceives the population as stagnated; dropping below it triggers a shift to an Exploration regime. The sensitivity analysis in Figure 4 reveals the delicate balance required here. A low threshold ($\delta_{p}=0.1$) proves overly tolerant of redundancy, allowing the LLM to repeatedly output identical text without intervention. This failure to trigger exploration traps the search in local optima, leading to premature convergence. Conversely, an excessively high threshold ($\delta_{p}=0.5$) imposes a hypersensitive constraint that misinterprets healthy convergence as stagnation. As evidenced by the significant instability in the green curve, this causes frequent, unnecessary disruptions that abort the fine-tuning of high-potential heuristics. Therefore, the default setting ($\delta_{p}=0.3$) represents a robust operating point. By triggering intervention only when over 70% of the population consists of exact duplicates, it functions as an indicator that effectively anticipates local-optimum collapse, enforcing proactive exploration before the search fully stalls while still permitting sufficient iterations for the local refinement of elite patterns.
Population Size ($N_{\text{pop}}$).
This parameter determines the breadth of the heuristic search space covered per generation. The sensitivity analysis in Figure 4 reveals a distinct trade-off: excessively small populations (e.g., $N_{\text{pop}}=4$) lead to premature convergence, as the limited gene pool lacks sufficient diversity to sustain effective recombination. Conversely, overly large populations (e.g., $N_{\text{pop}}=12$) significantly increase the computational burden (token consumption) without yielding proportional performance gains. The results confirm that a size of 8 serves as an effective balance point, providing sufficient genetic material for the LLM to synthesize high-quality heuristics while maintaining a manageable query budget.
Maximum Generations ($G_{\text{max}}$).
This parameter defines the temporal budget of the evolutionary process. The convergence trajectories in Figure 4 demonstrate that the algorithm typically achieves near-optimal solutions within the first 6–8 generations. While extending the search to 12 or more generations might offer slight refinements, it yields diminishing returns. The marginal performance improvements observed in these later stages do not justify the significant additional inference costs. Consequently, we selected $G_{\text{max}}=8$ as the standard configuration, representing a trade-off that secures high solution quality while maximizing computational efficiency.
Insight Pool Size ($C_{\text{pool}}$).
The capacity of the Insight Pool regulates the system’s long-term memory. As indicated by the variance and trends in Figure 4, a small pool (e.g., <10<10) leads to catastrophic forgetting, where valuable high-level principles are evicted before they can be effectively reused. On the other hand, an unconstrained or overly large pool acts as a noise accumulator, diluting high-quality insights with mediocre ones and confusing the LLM with irrelevant context. Our experiments indicate that a capacity of 30 effectively balances the retention of diverse, high-utility insights with the need to filter out noise, maintaining both stability and agility.
Stagnation Threshold ($\tau_{\text{stag}}$).
This parameter determines the sensitivity of the Evolutionary Navigator in detecting performance plateaus. It sets the number of consecutive non-improving generations required to trigger a shift from Exploitation to Exploration. The sensitivity analysis in Figure 4 highlights the critical balance involved. A low threshold (e.g., $\tau_{\text{stag}}=1$) makes the system hypersensitive, frequently interrupting the refinement of promising solutions due to minor stochastic fluctuations. This prevents the algorithm from fully exploiting local basins of attraction. Conversely, a high threshold (e.g., $\tau_{\text{stag}}=5$) renders the framework sluggish. Given the constrained evolutionary window (typically 12 generations) and the high cost of LLM inference, waiting for 5 generations represents a significant waste of the computational budget on unproductive search. The default value of $\tau_{\text{stag}}=3$ provides a trade-off, allowing for rapid detection of diminishing returns while filtering out transient noise in the optimization trajectory.
Figure 4: Parameter sensitivity analysis of population size, maximum generations, insight pool size, stagnation threshold and diversity threshold.
C.7 Additional Ablation Studies and Analysis
To provide a comprehensive understanding of the internal mechanisms of HiFo-Prompt and to dissociate the contribution of specific components from the overall framework, we conducted a series of additional ablation studies. These experiments investigate the architectural necessity of the Insight Pool, the efficacy of the adaptive Evolutionary Navigator, the behavioral dynamics of regime switching, and the fine-grained design of the utility function.
Architectural Necessity of the Insight Pool.
To determine whether the performance gains of HiFo-Prompt stem principally from the initialization with high-quality seed insights or from the framework’s capability to manage them, we injected the identical set of seed insights into the initial populations or prompts of four state-of-the-art baselines: EoH, ReEvo, HSEvo, and MCTS-AHD. As detailed in Table 13, the incorporation of seed insights into these baselines yields negligible performance improvements, and in certain cases (e.g., MCTS-AHD), results in slight performance degradation due to interference with their native search logic. This stands in sharp contrast to the significant gains achieved by HiFo-Prompt. These results substantiate that standard LLM-based evolutionary methods lack the necessary architectural mechanisms—specifically, the Insight Pool and the Hindsight reflection process—to effectively retain, retrieve, and utilize abstract knowledge. Consequently, the superior performance of HiFo-Prompt is attributed to its unique ability to leverage high-level guidance rather than merely the quality of initialization.
Table 13: Impact of seed insights on baselines, using TSP construction as an example. The results show the objective function values. Lower is better.
Method TSP10 TSP20 TSP50 TSP100 TSP200
EoH (w/ seed insight) 3.024 4.233 6.512 8.998 12.470
EoH (w/o seed insight) 3.049 4.252 6.441 8.990 12.514
ReEvo (w/ seed insight) 2.993 4.132 6.344 8.759 12.198
ReEvo (w/o seed insight) 2.994 4.126 6.294 8.773 12.324
HSEvo (w/ seed insight) 3.064 4.324 6.452 8.842 12.270
HSEvo (w/o seed insight) 3.001 4.170 6.307 8.729 12.183
MCTS-AHD (w/ seed insight) 2.938 4.142 6.294 8.724 12.076
MCTS-AHD (w/o seed insight) 2.983 4.174 6.317 8.769 12.176
Efficacy of Adaptive Regime Switching.
The Evolutionary Navigator is designed to dynamically switch between Explore, Exploit, and Balance regimes based on real-time population states. To validate the necessity of this adaptive control, we compared the full HiFo-Prompt against a variant fixed permanently in the Balance mode, which attempts to combine exploration and exploitation in every prompt without explicit directional guidance. Table 14 demonstrates that the adaptive mechanism significantly outperforms the fixed strategy. This performance gap widens as the problem scale increases (e.g., reducing the gap on TSP200 from 22.115% to 8.877%). This indicates that a static prompt strategy is insufficient for navigating complex search landscapes. The Navigator’s ability to explicitly enforce exploration during stagnation and exploitation during progress is critical for escaping local optima and refining solutions efficiently.
Table 14: Ablation study of the Evolutionary Navigator’s adaptive mechanism vs. a fixed strategy.
Method TSP10 TSP20 TSP50 TSP100 TSP200
HiFo-Prompt (adaptive) 1.654% 3.619% 6.625% 8.582% 8.877%
HiFo-Prompt (fixed Balance) 5.812% 9.676% 15.599% 18.924% 22.115%
Behavioral Analysis of the Evolutionary Navigator.
To further elucidate the operational behavior of the Navigator, we analyzed the frequency of each regime triggered during the evolutionary process across three independent runs. As shown in Table 15, the system exhibits a balanced distribution of states. While Balance is the most frequent state (providing stability), the system frequently triggers Explore (avg. 102.3 times) and Exploit (avg. 83.7 times). This confirms that the diversity ($\delta_{p}$) and stagnation ($\tau_{stag}$) thresholds are actively functioning, ensuring the search dynamically oscillates between structural diversification and local refinement rather than converging prematurely.
Table 15: Frequency statistics of Evolutionary Regimes triggered by the Navigator.
Regime run 1 run 2 run 3 avg.
Explore 110 95 102 102.3
Exploit 81 88 82 83.7
Balance 129 137 136 134.0
Synergistic Effect of Utility Function Components.
The utility function $U(k_{i},t)$ is not merely a sum of terms but establishes a dynamic interplay among three constituents to ensure stable, convergent selection. The learned effectiveness $E_{i}(t)$ serves as the dominant, low-variance signal reflecting long-term value via Exponential Moving Average (EMA). The usage penalty and recency bonus act as critical modulating factors. Specifically, the usage penalty employs a logarithmic growth design ($\log(N_{i}(t)+1)$); this ensures the penalty saturates quickly, meaning an insight’s priority is gently down-weighted rather than suffering a precipitous decline. This prevents sharp priority reversals and oscillatory behavior while preserving diversity. Conversely, the recency bonus injects a fixed, time-dependent reward to encourage search continuity along recently validated trajectories.
The ablation study in Table 16 empirically confirms this synergy. Removing the penalty ($w_{u}=0$) leads to inferior results, indicating that without exploration pressure, the system over-exploits early dominant insights, leading to premature knowledge collapse. Removing the reward ($\tau_{r}=0$) similarly harms performance, as the search loses its momentum and explores widely but inefficiently. The poorest performance occurs when both modulators are removed, validating that the simultaneous application of penalizing overuse and rewarding recency is essential for robust credit assignment.
Table 16: Ablation analysis of usage penalty ($w_{u}$) and recency bonus ($\tau_{r}$) in the Utility Function.
Setting run 1 run 2 run 3 avg.
$w_{u}>0,\tau_{r}>0$ (default) 5.885 5.838 5.873 5.866
$w_{u}=0,\tau_{r}>0$ (w/o penalty) 5.958 5.989 5.978 5.975
$w_{u}>0,\tau_{r}=0$ (w/o reward) 5.954 5.972 5.979 5.968
$w_{u}=0,\tau_{r}=0$ (w/o both) 5.984 6.026 5.982 5.997
Calibration of Base Credit Assignment.
Finally, we analyzed the sensitivity of the base credit assignment coefficients defined in Eq. 3 ($g_{\text{eff}}^{\text{raw}}$). In Table 17, each configuration is presented as a triplet $\beta_{\text{best}},\beta_{\text{inc}},\beta_{\text{pen}}]$, representing the base intercept values for three distinct performance tiers: the first value ($\beta_{\text{best}}$) rewards offspring that surpass the current population best (paradigm-shifting); the second ($\beta_{\text{inc}}$) rewards those improving upon the average but not the best (incremental); and the third ($\beta_{\text{pen}}$) penalizes individuals performing below the average (sub-optimal).
Table 17: Sensitivity analysis of base reward configurations.
Configuration run 1 run 2 run 3 avg.
$[0.8,0.2,0.0]$ 6.046 5.976 5.998 6.007
[1.0,0.1,0.5][1.0,0.1,-0.5] 5.921 5.932 5.902 5.918
[0.6,0.3,0.1][0.6,0.3,-0.1] 5.916 5.917 5.915 5.916
[0.8,0.2,0.3][0.8,0.2,-0.3] (default) 5.885 5.838 5.873 5.866
Our default configuration [0.8,0.2,0.3][0.8,0.2,-0.3] is not arbitrary but follows a specific asymmetric design philosophy: $\text{Elite Reward}\gg|\text{Penalty}|\gtrsim\text{Incremental Reward}$. The substantial elite reward (0.8) provides a strong signal for breakthroughs, prioritizing paradigm discovery. The incremental reward (0.2) acknowledges smaller gains, while the penalty (-0.3)—larger in magnitude than the incremental reward—effectively prunes unproductive directions. This asymmetry improves search efficiency by aggressively filtering noise while amplifying true innovations. Additionally, the coefficients follow a $\wedge$-shaped sensitivity design, where the intermediate tier receives the largest gradient to encourage rapid upward transitions, while the elite tier is assigned lower sensitivity to maintain stability. The results in Table 17 confirm that this theoretically grounded configuration yields the lowest objective value. Deviating towards more aggressive rewards (e.g., [1.0,0.1,0.5][1.0,0.1,-0.5]) or different penalty structures disrupts this balance, resulting in inferior performance.
Standard Deviation Analysis.
To comprehensively evaluate the stability and consistency of the models, we conducted three independent runs of all LLM-based AHD methods on several representative scales. We then summarized the objective function values of each run and reported their average (avg.) and standard deviation (std.), as shown in Table 18. These statistics provide a clear indication of the variability across runs, offering a more reliable basis for subsequent performance comparison and robustness analysis.
Table 18: Objective function values of three run and their average(avg.) and standard deviation(std.).
Task Method run1 run2 run3 avg. std.
TSP50-construction EoH 6.369 6.554 6.401 6.441 0.09887534
ReEvo 6.264 6.298 6.320 6.294 0.02821347
HSEvo 6.348 6.259 6.314 6.307 0.04491102
MCTS-AHD 6.359 6.288 6.303 6.317 0.03742103
Ours 6.135 6.087 6.041 6.088 0.04700355
TSP100-GLS EoH 7.738 7.738 7.739 7.738 0.00034451
ReEvo 7.742 7.741 7.738 7.740 0.00173957
HSEvo 7.739 7.737 7.753 7.743 0.00894281
Ours 7.737 7.737 7.738 7.737 0.00007013
Online BPP 5k_C100 EoH 2028 2027 2025 2027 1.11355287
ReEvo 2021 2023 2022 2022 0.60880932
HSEvo 2022 2024 2088 2045 37.59627641
MCTS-AHD 2027 2026 2026 2026 0.66901381
Ours 2019 2020 2021 2020 0.64291005
FSSP-GLS n100_m10 EoH 5640 5645 5644 5643 2.60832002
Ours 5635 5638 5638 5637 1.67470225
C.8 Consumption of Time and Token
Compared to other LLM-based AHD methods, our approach demonstrates advantages in both time and token consumption, with particularly significant reductions in computational time. Based on qwen2.5-max model, we calculated the runtime and token consumption of various methods across three problem domains: TSP-construction, Online BPP, and BO-CAF in Table 19.
Table 19: Time and token consumption with different methods.
Methods Consumption TSP-construction Online BPP BO-CAF
EoH Time 1.2h 1h 2h
Input Token 0.8M 0.5M 1.2M
Output Token 0.2M 0.2M 0.5M
ReEvo Time 2h - -
Input Token 1.1M - -
Output Token 0.4M - -
MCTS-AHD Time 4h 3h 14h
Input Token 1M 1M 1.3M
Output Token 0.3M 0.2M 0.6M
Ours Time 40 min 36min 1h
Input Token 0.5M 0.28M 0.8M
Output Token 0.2M 0.12M 0.2M
Appendix D Limitation and Future Work
D.1 Limitation